OpenStudy (anonymous):

will medal(: Which of these is a simplified form of the equation 7y + 7 = 9 + 2y + 2y?

OpenStudy (anonymous):

3y = 2 7y = 6 11y = 16 4y = 16

OpenStudy (anonymous):

please help(: @Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

A. Confirm it. Credit to: Openstudy

OpenStudy (kash_thesmartguy):

Did I help?

OpenStudy (kash_thesmartguy):

@leahh

OpenStudy (anonymous):

thank you so much! @Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

Okay, tell if you have any others!

OpenStudy (anonymous):

i do!

OpenStudy (anonymous):

do you mind to help me on a few more? @Kash_TheSmartGuy

OpenStudy (anonymous):

What is the value of y in the equation 3(3y − 12) = 0? 4 5 6 9

OpenStudy (pawanyadav):

Taking variables and constant on opposite sides of equal to you will get the answer A

OpenStudy (anonymous):

okay thanks! @Pawanyadav

OpenStudy (anonymous):

can you answer another

OpenStudy (pawanyadav):

9y-36=0 y=4

OpenStudy (anonymous):

What is the value of x in the equation −6 + x = −1?

OpenStudy (anonymous):

−5 −7 7 5

OpenStudy (kash_thesmartguy):

5

OpenStudy (anonymous):

thanks! @Kash_TheSmartGuy

OpenStudy (anonymous):

What is the solution to the equation 1.6m − 4.8 = −1.6m? m = 0.5 m = 0.7 m = 1.5 m = 3

OpenStudy (anonymous):

@Kash_TheSmartGuy @pamela.g

OpenStudy (anonymous):

@Pawanyadav

OpenStudy (pawanyadav):

3.2m=4.8 m=1.5

OpenStudy (kash_thesmartguy):

m=1.5 or D. Check out: http://www.geteasysolution.com/1.6m-4.8=-1.6m

OpenStudy (pawanyadav):

3.2m=4.8 m=1.5

OpenStudy (pawanyadav):

It can't be 3